jaelaj3:

The graph below shows the number of paintballs a machine launches, y, in x seconds: A graph titled Rate of Launch is shown. The x axis label is Time in seconds, and the x axis values are from 0 to 10 in increments of 2 for each grid line. The y axis label is Number of Balls, and the y axis values from 0 to 60 in increments of 12 for each grid line. A line is shown connecting points on ordered pair 2, 12 and 4, 24 and 6, 36 and 8, 48. Which expression can be used to calculate the rate per second at which the machine launches the balls?

jaelaj3:

I know the answer, I just need an explanation. please help

koenko:

what level is this

jaelaj3:

7th grade math

koenko:

oh okay

koenko:

we know that The rate of a linear equation is equal to the slope m The slope is equal to points (2,12) and (4,24) m=(24-12)/(4-2) m=12/2 m=6 balls/sec

jaelaj3:

its rise over run. so its 12 for every two seconds, 2/12, so basically 6 for every second. I just dont know how to explain this.

koenko:

im confused

jaelaj3:

never mind U figured it out. Thanks!

koenko:

np